# Dhaka 1229, Bangladesh

### v1

@@ Abstract
This paper analyzes the online facility assignment problem in a geometric setting where facilities with unit capacity are positioned at the vertices of a regular $n$-gon. Customers arrive sequentially at uniformly random positions along the edges. They must be assigned immediately to the nearest available facility, with ties broken by coin toss. The sequential nature and unknown future arrivals require a probabilistic analysis of the expected assignment cost. Our main contribution is a recursive characterization of the expected cost: for any occupancy state $S$, the expected remaining cost $V(S)$ equals the average over all edge positions of the immediate assignment cost plus the expected future cost after assignment. We prove that this integral equation can calculate a solution and provide the expected value for small $n$ ($n=3,4,5$). For larger values of $n$ and expected cost, we develop efficient numerical methods, including a discretized dynamic programming approach and Monte Carlo simulation. The work establishes a fundamental probabilistic approach for online assignment in polygonal environments.
@@ 1.3 Contributions
    where $C_{S,e}(t)$ is the immediate cost and $P_{m|S,e}(t)$ are transition probabilities.
@@
    Computational Characterization: Through experiments for $n=3$ to $12$, we provide expected costs and analyze scaling behaviour with polygon size.
@@ 2 Related Work
The problem of assigning customers to facilities in an online setting sits at the intersection of online algorithms, geometric optimization, and stochastic analysis. The classical Online Facility Location (OFL) problem, where the algorithm must decide whether to open a facility or connect to an existing one, was formalized by Meyerson [11]. He provided a randomized algorithm with a logarithmic competitive ratio, a result that has been refined and generalized for various metric spaces [4]. However, our problem differs from OFL in that facility locations are fixed and capacities are strict. This variant, formally known as the Online Facility Assignment (OFA) problem, has been studied by Ahmed et al. [1], who proved that a greedy strategy yields a linear competitive ratio on line and graph geometries. Recently, Malik et al. [8] extended this to regular polygon geometries, focusing on worst-case bounds.
Our work complements these competitive analyses by addressing the average-case scenario, which shares theoretical roots with the Online Bipartite Matching problem. The seminal work by Karp et al. [7] established that for adversarial arrivals, the optimal competitive ratio is 11/e1-1/e. However, when the assumption of adversarial arrivals is relaxed to random order arrivals (ROA), algorithms often perform significantly better. Goel and Mehta [5] demonstrated that the competitive ratio approaches 1 as the size of the instance grows under random permutations. For weighted variants involving distances, Kalyanasundaram and Pruhs [6] showed that simple greedy algorithms can fail in adversarial settings but perform well stochastically. A comprehensive survey of these online matching techniques and their applications in resource allocation is provided by Mehta [9].
In the specific context of geometric probability, the expected cost of matching random points has a deep theoretical concept. The result discussed by Ajtai, Komlós, and Tusnády [2] says that for $n$ random points in a unit square, the optimal matching cost scales as $\Theta(\sqrt{n\log n})$. However, their work provides asymptotic bounds for large $n$ and does not offer methods for calculating exact costs for finite $n$. Furthermore, as the state space in these problems grows exponentially, the determination of exact analytical solutions becomes computationally intractable. To address this, Monte Carlo methods [10] are frequently employed to approximate high-dimensional integrals and expectations in stochastic systems. Caflisch [3] highlights that for problems where standard numerical integration fails due to dimensionality, Monte Carlo convergence rates remain independent of dimension. Our work synthesizes these approaches: we rely on the recursive structure for exactness in small instances and leverage Monte Carlo integration to estimate the expected costs for larger polygon sizes.
@@ 3 Problem Formulation
Figure 1: Distance Calculation Method
Customers arrive according to a uniform stochastic process: each customer independently selects an edge $e\in E$ with probability $\frac{1}{n}$, then a position $t\in[0,1]$ uniformly along that edge. Here $t=0$ corresponds to the starting vertex of the edge and $t=1$ to the ending vertex.
@@
$d(e,t,0)$ $\implies\min\left(t+\delta(F_{3},F_{0}),\ (1-t)+\delta(F_{2},F_{0})\right)$
$\implies\min\left(t+2,\ (1-t)+2\right)$
$\implies t+2$
@@ 4.1 Expected Cost Recurrence
$V(S)=\frac{1}{n}\sum_{e\in E}\int_{0}^{1}\left[C_{S,e}(t)+\sum_{m=1}^{n}P_{m|S,e}(t)\cdot V(S\cup\{m\})\right]dt$
@@ 5.1 Worked example: Exact expected cost for a square ($n=4$)
$\begin{array}[]{ll}\text{Edge }(0,1):&C(t)=1-t,\quad\int_{0}^{1}(1-t)\,dt=\tfrac{1}{2},\\[2.0pt] \text{Edge }(1,2):&C(t)=\min(t,1-t),\quad\int_{0}^{1}\min(t,1-t)\,dt=\tfrac{1}{4},\\[2.0pt] \text{Edge }(2,3):&C(t)=\min(t,1-t),\quad\int_{0}^{1}\min(t,1-t)\,dt=\tfrac{1}{4},\\[2.0pt] \text{Edge }(3,0):&C(t)=t,\quad\int_{0}^{1}t\,dt=\tfrac{1}{2}.\end{array}$

### v2

@@ Abstract
This paper analyzes the online facility assignment problem in a geometric setting where facilities with unit capacity are positioned at the vertices of a regular $n$-gon. Customers arrive sequentially at uniformly random positions along the edges. They must be assigned immediately to the nearest available facility, with ties broken by coin toss. The sequential nature and unknown future arrivals require a probabilistic analysis of the expected assignment cost. Our main contribution is a recursive characterization of the expected cost: for any occupancy state $S$, the expected remaining cost $V(S)$ equals the average over all edge positions of the immediate assignment cost plus the expected future cost after assignment. We prove that this integral equation can calculate a solution and provide the expected value for small $n$ ($n=3,4,5,\dots$). For larger values of $n$ and expected cost, we develop efficient numerical methods, including a discretized dynamic programming approach and Monte Carlo simulation. The work establishes a fundamental probabilistic approach for online assignment in polygonal environments.
@@ 1.1 Problem Motivation
Beyond modeling these systems, our expected-cost analysis can provide practical value for real-world decision-making. By understanding the average distance a customer must travel before being served, companies can derive actionable insights such as:
  1. i.
    Predicting resource requirements: Organizations can determine the number of drivers, delivery vehicles, or service nodes required to maintain a desired service level.
  2. ii.
    Optimizing facility placement and routing: Even under online or uncertain demand, expected-cost trends help identify configurations that minimize travel distance.
  3. iii.
    Reducing operational expenses: Lower expected distances imply reduced fuel consumption, travel time, and congestion.
  4. iv.
    Improving customer experience: Companies can use expected cost to guarantee faster response times and formulate service-level agreements.
  5. v.
    Supporting scalable planning: Since expected cost grows predictably with the number of facilities, organizations can anticipate performance as their systems expand.
Thus, analyzing the expected assignment cost is not only theoretically significant but also practically valuable, helping companies design robust, cost-effective, and customer-friendly systems operating under uncertainty.
@@ 1.3 Contributions
    Where:
    1. i.
      $S$ = the set of already occupied facilities.
    2. ii.
      $e$ = an edge chosen uniformly from the $n$ edges.
    3. iii.
      $t\in[0,1]$ = the position of arrival on that edge.
    4. iv.
      $C_{S,e}(t)$ = the immediate cost of connecting the arrival to the nearest free facility.
    5. v.
      $P_{m\mid S,e}(t)$ = the probability that facility $m$ becomes the next occupied one (depends on distance and tie-breaking).
    6. vi.
      $V(S)$ = the expected future cost starting from state $S$.
@@
    Computational Characterization: Through experiments for $n=3$ to $20$, we provide some expected costs and analyze scaling behaviour with polygon size.
@@ 2 Related Work
The problem of assigning customers to facilities in an online setting sits at the intersection of online algorithms, geometric optimization, and stochastic analysis. The classical Online Facility Location (OFL) problem, where the algorithm must decide whether to open a facility or connect to an existing one, was formalized by Meyerson [meyerson2001online]. He provided a randomized algorithm with a logarithmic competitive ratio, a result that has been refined and generalized for various metric spaces [fotakis2011online]. However, our problem differs from OFL in that facility locations are fixed and capacities are strict. This variant, formally known as the Online Facility Assignment (OFA) problem, has been studied by Ahmed et al. [ahmed2020online], who proved that a greedy strategy yields a linear competitive ratio on line and graph geometries. Recently, Malik et al. [malik2025online] extended this to regular polygon geometries, focusing on worst-case bounds.
Our work complements these competitive analyses by addressing the average-case scenario, which shares theoretical roots with the Online Bipartite Matching problem. The seminal work by Karp et al. [karp1990optimal] established that for adversarial arrivals, the optimal competitive ratio is 11/e1-1/e. However, when the assumption of adversarial arrivals is relaxed to random order arrivals (ROA), algorithms often perform significantly better. Goel and Mehta [goel2008online] demonstrated that the competitive ratio approaches 1 as the size of the instance grows under random permutations. For weighted variants involving distances, Kalyanasundaram and Pruhs [kalyanasundaram1993online] showed that simple greedy algorithms can fail in adversarial settings but perform well stochastically. A comprehensive survey of these online matching techniques and their applications in resource allocation is provided by Mehta [mehta2013online].
In the specific context of geometric probability, the expected cost of matching random points has a deep theoretical concept. The result discussed by Ajtai, Komlós, and Tusnády [ajtai1984optimal] says that for $n$ random points in a unit square, the optimal matching cost scales as $\Theta(\sqrt{n\log n})$. However, their work provides asymptotic bounds for large $n$ and does not offer methods for calculating exact costs for finite $n$. Furthermore, as the state space in these problems grows exponentially, the determination of exact analytical solutions becomes computationally intractable. To address this, Monte Carlo methods [metropolis1949monte] are frequently employed to approximate high-dimensional integrals and expectations in stochastic systems. Caflisch [caflisch1998monte] highlights that for problems where standard numerical integration fails due to dimensionality, Monte Carlo convergence rates remain independent of dimension. Our work synthesizes these approaches: we rely on the recursive structure for exactness in small instances and leverage Monte Carlo integration to estimate the expected costs for larger polygon sizes.
@@ 3 Problem Formulation
Customers arrive according to a uniform process: each customer independently selects an edge $e\in E$ with probability $\frac{1}{n}$, then a position $t\in[0,1]$ uniformly along that edge. Here $t=0$ corresponds to the starting vertex of the edge and $t=1$ to the ending vertex.
@@
$\implies d(e,t,0)$ $=\min\left(t+\delta(F_{3},F_{0}),\ (1-t)+\delta(F_{2},F_{0})\right)$
$\implies d(e,t,0)$ $=\min\left(t+2,\ (1-t)+2\right)$
$\implies d(e,t,0)$ $=t+2$
@@ 4.1 Expected Cost Recurrence
Component Meaning Why Used
$S$ Current occupancy set Describes the current state of which facilities are already occupied.
$V(S)$ Expected future cost from state $S$ Objective function; required to compute total expected cost through dynamic programming.
$\frac{1}{n}$ Uniform distribution over $n$ edges Customers arrive uniformly across edges, so expectation requires dividing by $n$.
$\sum_{e\in E}$ Summation over all edges Arrival can occur on any edge; we must include all possibilities.
$\int_{0}^{1}dt$ Continuous arrival position on edge Arrival position is random along the edge; integration takes expectation over position.
$C_{S,e}(t)$ Immediate cost of connecting arrival to nearest free facility Measures the cost incurred for this specific arrival before future arrivals occur.
$m\notin S$ unoccupied facilities Only free facilities can be assigned, so they determine assignment options.
$P_{m\mid S,e}(t)$ Probability that facility $m$ becomes the next occupied facility Used for randomized decisions or ties where multiple nearest facilities exist.
$V(S\cup\{m\})$ Expected cost after occupying facility $m$ Captures how the future changes after the current assignment.
$\sum_{m\notin S}$ Expectation over the next possible facility to be occupied Computes expected future cost over all assignment outcomes.
Table 1: Explanation of all components used in the expected cost equation.
@@
$V(S)=\frac{1}{n}\sum_{e\in E}\int_{0}^{1}\left[C_{S,e}(t)+\sum_{m\notin S}^{n}P_{m|S,e}(t)\cdot V(S\cup\{m\})\right]dt$
@@ 5.1 Worked example: Exact expected cost for a square ($n=4$)
$\begin{array}[]{ll}\text{Edge }(0,1):&C(t)=1-t,\implies\int_{0}^{1}(1-t)\,dt=\tfrac{1}{2},\\[2.0pt] \text{Edge }(1,2):&C(t)=\min(t,1-t),\implies\int_{0}^{1}\min(t,1-t)\,dt=\tfrac{1}{4},\\[2.0pt] \text{Edge }(2,3):&C(t)=\min(t,1-t),\implies\int_{0}^{1}\min(t,1-t)\,dt=\tfrac{1}{4},\\[2.0pt] \text{Edge }(3,0):&C(t)=t,\implies\int_{0}^{1}t\,dt=\tfrac{1}{2}.\end{array}$